### v1

# A General Bayesian Framework for Informative Input Design in System Identification

Abstract
We tackle the problem of informative input design for system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data. We propose a methodology that is compatible with any system and parametric family of models. Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters. Our algorithm consists of two modules. First, we formulate the problem of system identification from a Bayesian perspective and propose an approximate iterative method to optimize the model’s parameters. Based on this Bayesian formulation, we are able to define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input. Our method outperforms model-free baselines with various linear and nonlinear dynamics.
I Introduction
Controlling uncertainty is a central problem for various engineering fields. It abstracts to the selection of inputs to a system in order to reduce an uncertainty measure for a variable of interest. In robotics, the problem usually coincides with adaptive informative input design, where trajectories are planned to maximize the information gathered from the unknown environment [1]. Similarly, robots may choose actions to reduce the uncertainty in their state [2]. In systems engineering, we may want to choose inputs to a system to improve the estimate of system dynamics [3].
In this work, we are interested in the problem of informative input design for the modeling of a system. Our formulation tackles the general problem of dataset collection for model-fitting purposes. Ott et al. have shown that selecting control inputs based on the current model can improve the system identification process [4]. Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications. This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging. Informative input design can also be regarded as the first step in a more complicated pipeline that involves the control of the underlying system, e.g., with model predictive control [5]. The problem of selecting input sequences that lead to highly informative measurements under constraints is still an open problem [4].
There are four basic elements in the problem of informative input design. We need to select a parametric family of models for the underlying system. Our parameter estimate is improved once more data becomes available and we must choose the optimization approach to update the model parameters. The model family must be expressive enough to capture the true system, but also simple enough to allow the optimization algorithm to converge in a meaningful manner. Furthermore, we need an uncertainty measure to minimize over the next input. This allows us to select inputs that best inform our choice of model parameters. The last minimization problem must also be accompanied by an optimization approach.
These specifications make designing methods for informative input design challenging and lead to algorithms that are usually limited to a specific set of parametric models. Our approach allows for the selection of any parametric family of models for the true system, although we do not focus on the suitability of the chosen model. It leverages first-order information for the model and successive convex problems based on the first-order Taylor expansion to update the model parameters. This simplifies the model update step discussed in the previous paragraph. Because the first-order Taylor expansion gives rise to a linear Gaussian model, we use an uncertainty measure that is accurate in the linear Gaussian case for the input design.
The rest of the paper is organized as follows. In section II we discuss prior work. In section III, we formulate the problem of informative input design, while in section IV we describe our proposed approach. We test our algorithm in various scenarios in section V and the paper concludes with section VI.
II Prior Work
The problem of informative input design has been well studied. Early research on the optimal selection of inputs dates back to the 1960s [6, 7] and highlights the importance of the problem for the identification of systems. Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [8, 9]. This section summarizes three categories of informative input design: model-free methods, methods for linear systems, and methods for nonlinear systems.
Informative inputs can be designed by ignoring the current estimate of the system dynamics. These methods are called model-free because they are not based on a model of the system. One such method is the orthogonal optimized multisine input design [10], which has found applications in the study of aircraft stability and control. This technique produces inputs that are mutually orthogonal in both the time and frequency domains, in an attempt to perturb the system in orthogonal directions. Similarly, pseudorandom binary sequences can be used to excite the system [11]. The idea is to leverage the sequence’s randomness in order to excite the system in interesting ways. Our method differs from model-free approaches because it uses the currently estimated model to inform the next input design.
Although they are simple, linear models are a powerful representation tool. They usually allow for closed-form analysis and are easily interpretable from a statistical standpoint. Informative input design using such models has a vast literature. A recent method that aims to estimate linear dynamics suggests selecting the next input by minimizing the trace of the covariance of the model parameters with linear dynamics constraints [4]. By linearizing the objective, the authors obtain a convex problem that is easily solvable for the next input. This method is compatible with the dynamic mode decomposition with control (DMDc) framework that produces reduced order models for the dynamics [12]. Our method can provide an approximate solution to any parametric model family and is exact in the case of linear models.
In various applications, the dynamics contain modes that are not properly described by a linear model. In these cases, we must design inputs for nonlinear system identification. Early results focused on specific dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14]. In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification. In this case, convex-based methods exist for informative input design [15]. In contrast with these methods, our proposed algorithm is applicable in the case of general nonlinear models.
Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16]. Our work adds to this work in an important way: we present a method to update the estimate of the model parameters that respects the linearization and is based on convex optimization, which makes it computationally fast. A similar work was done by Babar and Baglietto [17], who further propose an extended Kalman filter for the parameter estimation. We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input. Our objective depends on a parameter that directly encourages the satisfaction of any constraints, i.e., dynamics constraints. In contrast with both works [16, 17], we present a fine-tuning method for the assumed noise covariances that can improve convergence.
III Problem Formulation
We assume that a system $g$ is characterized by the relation
$y=g(x),$ (1)
where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$. If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$. We refer to $x$ as the input and $y$ as the output.
We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{q}$. At every iteration, our goal is to compute:
  1. 1.
    $\hat{\theta}$ such that $g(\cdot)\approx f(\cdot;\hat{\theta})$ using the available data,
  2. 2.
    the optimal next input to the system that decreases our uncertainty over the unknown parameter $\theta$.
We opt for a Bayesian treatment of the problem. We suppose that the inputs and outputs are exactly observed, because our main goal is efficient dataset acquisition. Our method easily extends to the case of noisy observation of the outputs.
IV Proposed Method
We present our methodology for the design of an informative input in system identification. We assume that we have already observed input-output data points $(x_{i},y_{i})_{i=1}^{n}$ from the dynamics:
$y_{i}=g(x_{i}).$ (2)
We first outline our Bayesian approach for improving our current estimate of the model parameters. We then use this Bayesian formulation in order to define an optimization problem for the next input. Finally, we discuss some fine-tuning that can improve performance before bringing the concepts together in our proposed algorithm.
IV-A Parameter Estimation
The collected data satisfy eq. 2. Because the current model $f(\cdot;\theta)$ might not match the true system $g(\cdot)$ exactly, it holds that
$y_{i}=f(x_{i};\theta)+\epsilon_{i},$ (3)
for some noise realization $\epsilon_{i}$. We choose the underlying noise model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma_{i})$. Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
$p(\theta\mid y_{1},\dots,y_{n})\propto p(\theta)\prod_{i=1}^{n}p(y_{i}\mid\theta),$ (4)
where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$. The noise and prior are assumed Gaussian because they lead to a convex maximum likelihood problem [18].
The maximum a posteriori estimate for the parameter under these assumptions is given by
$\hat{\theta}_{+}$ $=\operatorname*{arg\;max}_{\theta}\log p(\theta\mid y_{1},\dots,y_{n})$ (5)
$=\operatorname*{arg\;min}_{\theta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}+$
$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma_{i}^{-1}}^{2}.$
In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$. We opt for an approximate treatment of the optimization problem that lends itself to a convex solver.
Suppose that the current parameter estimate is $\hat{\theta}$. We do not allow our next estimate to differ much relative to the current one. This is justified because the next input-output pair should not change our estimate drastically. For model parameters close to $\hat{\theta}$, the parametric family can be accurately approximated using a first order Taylor series. We write
$f(x;\theta)$ $\approx f(x;\hat{\theta})+\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\left(\theta-\hat{\theta}\right)$ (6)
$=b(x;\hat{\theta})+C(x;\hat{\theta})\theta$
for $\theta$ close to $\hat{\theta}$, where
$\left[\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\right]_{i,j}$ $=\dfrac{\partial f_{i}}{\partial\theta_{j}}|_{(x;\hat{\theta})},$ (7)
$b(x;\hat{\theta})$ $=f(x;\hat{\theta})-\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\hat{\theta},$
$C(x;\hat{\theta})$ $=\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}.$
Plugging (6) into (5) and imposing the distance constraint, we get that
$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta\lVert\hat{\theta}\rVert}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x;\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma_{i}^{-1}}^{2}.$
This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19]. The constraint guarantees that we search for the next model parameter in the region where the current Taylor series is accurate. It corresponds to a trust region.
IV-B Informative Input Design
After obtaining $\hat{\theta}_{+}$, our goal is to select $x_{n+1}$. The goal of selecting $x_{n+1}$ is to reduce our uncertainty in the posterior estimate of $\theta$. Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under completely linear Gaussian assumptions, the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
It has been shown that such an expression can be accurate in the nonlinear case, although it is exact only in the linear regime [20].
To minimize the posterior covariance, we maximize some notion of magnitude,
$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
of the posterior inverse covariance over $x_{n+1}$. We can do so because the uncertainty in the parameter estimate is proportional to the magnitude of the posterior covariance, which is inversely proportional to the magnitude of the inverse posterior covariance. For example, we can use $\log\det$ as $\mathcal{M}$. The determinant of the covariance equals the product of its eigenvalues, which are proportionally related to the semi-axes of the probability contours for a Gaussian distribution.
In the process of choosing $x_{n+1}$, there usually exist constraints. For example, the input might need to belong to a specific set, $x_{n+1}\in\mathcal{X}$. The same might hold for the corresponding output, $g(x_{n+1})\in\mathcal{Y}$. We can take such constraints into account by solving the following optimization problem:
$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
$+\lambda\left(\mathbbm{1}(x_{n+1}\in\mathcal{X})+\mathbbm{1}(f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y})\right).$
In order to solve eq. 11 with gradient descent methods, we can replace the indicators with smooth functions. By increasing parameter $\lambda\geq 0$, we can move the search towards feasible inputs $x_{n+1}$. This is an example of a penalty method [21]. As a receding horizon extension, we can easily extend eq. 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$.
Equation 9 defines the posterior covariance for the parameter $\theta$. At the next iteration , we can replace the constraint of eq. 8 with
$\lVert\theta-\hat{\theta}\rVert^{2}_{\Sigma_{\mathrm{posterior}}^{-1}}\leq\chi_{d,p}^{2},$
where $\chi_{d,p}^{2}$ is the value of the chi-squared distribution of $d$ degrees of freedom at the $p$-percentile. This would allow $\theta$ to vary within the contours that correspond up to the $p$-th percentile of the posterior Gaussian distribution. Choosing the adequate constraint in eq. 8 involves balancing this statistical interpretation with the accuracy of the Taylor series.
IV-C Updating the Covariances $\Sigma_{i}$
At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$. After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\hat{\epsilon}_{i}$ by comparing the true output to the expected output, i.e., for $i=1,\dots,n$
$\hat{\epsilon}_{i}=y_{i}-f(x_{i};\hat{\theta}_{+}).$ (12)
Using these samples, we can compute a sample covariance for the overall error in our model
$\hat{\Sigma}_{+}=w\Sigma_{\mathrm{meas,prior}}+\dfrac{1-w}{n}\sum_{t=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top},\ w\in\left[0,1\right],$ (13)
i.e., a convex combination of the empirical covariance and a prior covariance. This can be utilized in later iterations.
IV-D Algorithm
We combine the modules presented above in algorithm 1. We assume we have already acquired dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ and show the proposed procedure for augmenting the dataset by one point.
Notice that when computing $x_{n+1}$ in eq. 11 we choose $\Sigma_{i}\leftarrow\hat{\Sigma}_{+}$ from (13) for $i=1,\dots,n+1$. When performing the maximum a posteriori estimation in eq. 8 , however, we use $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. This selection is made because our next input should excite the system in a way that is informative and decreases the uncertainty present in our current system estimate. However, when improving the estimate of the model parameters we should use less informative covariances to robustify the estimation.
Algorithm 1 Proposed Algorithm for Parameter Estimation and Input Design
1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma_{\mathrm{meas,prior}}$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta,\lambda$ and $w$
2:$\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}},\ \mathrm{for\ all}\ i=1,\dots,n$
3:Obtain updated estimate $\hat{\theta}_{+}$ by eq. 8
4:$\Sigma_{i}\leftarrow\hat{\Sigma}_{+},\ \mathrm{for\ all}\ i=1,\dots,n+1$
5:Obtain next input $x_{n+1}$ by solving eq. 11
6:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
7:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$
V Results
We test our algorithm on both linear and nonlinear systems. [^1] The code is included in the repository: https://github.com/alextzik/informative_input_design We further perform an ablation study to show the importance of finetuning the covariances $\Sigma_{i}$ in our algorithm.
Baselines
Our emphasis is on the informative input design. We compare against three different model-free baselines for the selection of inputs. We maintain the same parameter estimation module from section IV in all methods. The first baseline is random. At every iteration, we run algorithm 1, with a change in line 4: we suppose $x_{n+1}\sim\mathcal{N}(0,I)$, i.e., $x_{n+1}$ is randomly selected. The second baseline is based on pseudorandom binary sequences and is called PRBS. At every iteration, we run algorithm 1, with a change in line 4: we suppose that each coordinate of $x_{n+1}$ originates from a pseudorandom binary sequence. Finally, the last baseline is called mutlisine and is based on orthogonal multisine inputs [10]. In line 4 of algorithm 1, we set the coordinate $i$ of $x_{n+1}$ as follows
$x_{n+1,i}=\sum_{j=0}^{M-1}A_{i,j}\sin\left(\dfrac{2\pi(d_{x}j+i)(n+1)}{T}+\phi_{i,j}\right),$
where $M$ is the total number of sinusoidal components and $T$ is the total number of inputs to be designed. We sample each phase $\phi_{i,j}$ uniformly in $\left[0,2\pi\right]$ and each amplitude $A_{i,j}$ uniformly in $\left[-1,1\right]$. We select $M=30$ in all experiments below.
Algorithm Parameters
We set $w=0.2$, $\delta=0.3$, and $M=30$. We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition. We do set an upper bound on the allowed input magnitude. If the input provided by any method has a larger magnitude, it is scaled to have the largest allowed magnitude. The first datapoint is assumed fixed for all methods.
Metrics
In all cases below , the true system belongs to the parametric model family. We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number. Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq. 9. We assume $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$. A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs. The results are gathered in table I for all experiments.
V-A Linear System
We first consider a linear system
$g(x_{t})=\begin{bmatrix}1&2\\ 3&4\end{bmatrix}x_{t}$ (14)
and the parametric model
$f(x_{t};\theta)=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{3}&\theta_{4}\end{bmatrix}x_{t}.$ (15)
Our results are shown in Figure 2. In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses. We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional. We further plot the designed inputs for the various methods. Our algorithm exhibits a faster convergence to the true model parameters compared to the baselines. This is verified by the lower value of $\log\det$ for the final error covariance.
V-B Hénon map dynamics
We consider the Hénon map dynamics
$g(x_{t})=\begin{bmatrix}1-\alpha x_{t,1}^{2}+x_{t,2}\\ \beta x_{t,1}\end{bmatrix},$ (16)
where $\alpha=1.4$ and $\beta=0.3$. It is one of the most studied systems with chaotic behavior [22]. The parametric model is
$f(x_{t};\theta)=\begin{bmatrix}1-\theta_{1}x_{t,1}^{2}+x_{t,2}\\ \theta_{2}x_{t,1}\end{bmatrix}.$ (17)
Our results are shown in Figure 3. Our proposed algorithm is able to converge to the true model parameters and outperforms the baselines with respect to both metrics considered. From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter the estimate has a higher uncertainty in the second coordinate $\theta_{2}$, which is captured by the elongated ellipse axes along that direction.
V-C Unicycle Model
We also evaluate our algorithm on the unicycle dynamics model, which is nonlinear and given by the equation
$x_{t+1}=\begin{bmatrix}x_{t,1}+u_{1,t}\Delta t\cos(x_{t,3})\\ x_{t,2}+u_{1,t}\Delta t\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\Delta t\end{bmatrix},$ (18)
where $u_{1,t}$ is the linear velocity input and $u_{2,t}$ is the angular velocity input. We use $\Delta t=0.1$. In this example, we will only be able to choose the control input $u_{t}$ at every timestep. In other words, the informative input design must be done in sequence, with the next state depending on the previous state and input. The parametric model is assumed to be
$f(x_{t},u_{t};\theta)=\begin{bmatrix}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}.$ (19)
The time discretization is thus one of the unknown parameters. We start the model from the origin with zero being the first control for all methods. Figure 4 shows the performance of our algorithm. It outperforms the considered baselines. We further plot $(x_{t,1},x_{t,2})$ for the various methods.
V-D Ablation
In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with a variant that uses $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$ throughout. The new variant is termed proposed (no finetuning). In this simulation , we remove the amplitude constraint. It is evident in Figure 1 that the variant does not converge properly. This highlights the importance of finetuning the covariances $\Sigma_{i}$.
Figure 1: Ablation study in the case of Hénon dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
Figure 2: Results for the linear dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) Inputs as a function of timestep for the various methods.
Figure 3: Results for the Hénon dynamics.
(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
(b) Model parameter estimate error over timesteps.
(c) State of dynamics as a function of timestep for the various methods.
Figure 4: Results for the unicycle dynamics.
Method/Experiment Linear Dynamics Hénon Dynamics Unicycle Dynamics
proposed $\mathbf{-18.5}$ $\mathbf{-14.9}$ $\mathbf{-9.8}$
random $-10.7$ $-5.9$ $-7.0$
PRBS $-10.2$ $-5.1$ $-9.2$
multisine $-13.7$ $-9.7$ $-8.9$
TABLE I: $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep for the various dynamics studied.
VI Conclusion
We present a Bayesian methodology for informative input design that is based on linear Gaussian assumptions. It extends to general nonlinear system identification through the Taylor expansion. A key module in our algorithm is the online calibration of the noise covariances that greatly impacts performance. Our algorithm also has limitations. We must select the trust region shape and volume when performing the parameter estimation , and we do not address the problem of selecting the parametric model family. Future work will propose a method for judging the suitability of a given parametric model family.
Acknowledgements
Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity. The NASA University Leadership Initiative (grant ##\##80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity. For the first author, this work was also partially funded through the Alexander S. Onassis Foundation Scholarship program. We would like to thank Joshua Ott and Arec Jamgochian for their feedback.

### v2

@@
-A General Bayesian Framework for Informative Input Design in System Identification
+An Iterative Bayesian Approach for System Identification based on Linear Gaussian Models
 Abstract
-We tackle the problem of informative input design for system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data.
-We propose a methodology that is compatible with any system and parametric family of models.
+We tackle the problem of system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data.
+We propose a flexible and computationally tractable methodology that is compatible with any system and parametric family of models.
 Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters.
 Our algorithm consists of two modules.
-First, we formulate the problem of system identification from a Bayesian perspective and propose an approximate iterative method to optimize the model’s parameters.
-Based on this Bayesian formulation, we are able to define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input.
-Our method outperforms model-free baselines with various linear and nonlinear dynamics.
+First, we formulate the problem of system identification from a Bayesian perspective and use a linear Gaussian model approximation to iteratively optimize the model’s parameters.
+In each iteration, we propose to use the input-output data to tune the covariance of the linear Gaussian model.
+This statistically calibrates the approach.
+Secondly, we define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input.
+We test our method with linear and nonlinear dynamics.
 I Introduction
-Controlling uncertainty is a central problem for various engineering fields.
-It abstracts to the selection of inputs to a system in order to reduce an uncertainty measure for a variable of interest.
-In robotics, the problem usually coincides with adaptive informative input design, where trajectories are planned to maximize the information gathered from the unknown environment [1].
-Similarly, robots may choose actions to reduce the uncertainty in their state [2].
-In systems engineering, we may want to choose inputs to a system to improve the estimate of system dynamics [3].
-In this work, we are interested in the problem of informative input design for the modeling of a system.
-Our formulation tackles the general problem of dataset collection for model-fitting purposes.
-Ott et al.
+In this work, we are interested in the problem of fitting a model to a system and choosing informative inputs to improve the model.
+Our formulation tackles the general problem of model-fitting and dataset collection, which is known as system identification.
+Obtaining a good model of the system is a key requirement in various engineering tasks.
+System identification is usually the first step in a larger pipeline that involves the control of the underlying system, e.g., with model predictive control [1].
+Fitting model parameters to data can be challenging when the model is a complicated function of the parameters.
+In addition, Ott et al.
 have shown that selecting control inputs based on the current model can improve the system identification process [2].
@@
 This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging.
-Informative input design can also be regarded as the first step in a more complicated pipeline that involves the control of the underlying system, e.g., with model predictive control [5].
-The problem of selecting input sequences that lead to highly informative measurements under constraints is still an open problem [4].
-There are four basic elements in the problem of informative input design.
+There are four basic elements in the problem of system identification.
 We need to select a parametric family of models for the underlying system.
-Our parameter estimate is improved once more data becomes available and we must choose the optimization approach to update the model parameters.
-The model family must be expressive enough to capture the true system, but also simple enough to allow the optimization algorithm to converge in a meaningful manner.
+The family must be expressive enough to capture the true system, but also simple enough to optimize over.
+Our parameter estimate is improved once more data becomes available and we must also choose the optimization approach that updates the model parameters.
+A vast literature exists on methods for estimating the model parameters [3].
 Furthermore, we need an uncertainty measure to minimize over the next input.
 This allows us to select inputs that best inform our choice of model parameters.
-The last minimization problem must also be accompanied by an optimization approach.
-These specifications make designing methods for informative input design challenging and lead to algorithms that are usually limited to a specific set of parametric models.
-Our approach allows for the selection of any parametric family of models for the true system, although we do not focus on the suitability of the chosen model.
-It leverages first-order information for the model and successive convex problems based on the first-order Taylor expansion to update the model parameters.
-This simplifies the model update step discussed in the previous paragraph.
-Because the first-order Taylor expansion gives rise to a linear Gaussian model, we use an uncertainty measure that is accurate in the linear Gaussian case for the input design.
+Various objectives based on the Fisher information matrix can be minimized over the input [4].
+This optimization problem must also be accompanied by an optimization approach.
+The specifications discussed above make designing methods for system identification challenging and lead to algorithms that are usually limited to a specific set of parametric models.
+Our work deals with the following question:
+  • Can we design a method for system identification that is computationally tractable, compatible with any parametric model family and able to suggest whether a parametric model family is suitable for a given system?
+Our approach leverages first-order information to construct a linear Gaussian model for the output given the input, as a function of the model parameters.
+Based on the first-order Taylor expansion, we solve successive convex problems, corresponding to maximum likelihood estimation (MLE), to update the model parameters.
+This simplifies the model update step discussed in the previous paragraph, which becomes computationally tractable.
+We also propose a procedure to tune the covariance associated with the linear Gaussian model.
+This allows us to statistically calibrate our model and judge the suitability of the parametric model family.
+Finally, using the linear Gaussian model from the first-order Taylor expansion, we define an uncertainty measure that is accurate in the linear Gaussian case for the input design.
 The rest of the paper is organized as follows.
 In section II we discuss prior work.
-In section III, we formulate the problem of informative input design, while in section IV we describe our proposed approach.
+In section III, we formulate the problem of system identification, while in section IV we describe our proposed approach.
 We test our algorithm in various scenarios in section V and the paper concludes with section VI.
 II Prior Work
-The problem of informative input design has been well studied.
+The problem of system identification has been well studied.
 Early research on the optimal selection of inputs dates back to the 1960s [5, 6] and highlights the importance of the problem for the identification of systems.
 Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [7, 8].
-This section summarizes three categories of informative input design: model-free methods, methods for linear systems, and methods for nonlinear systems.
-Informative inputs can be designed by ignoring the current estimate of the system dynamics.
-These methods are called model-free because they are not based on a model of the system.
-One such method is the orthogonal optimized multisine input design [10], which has found applications in the study of aircraft stability and control.
-This technique produces inputs that are mutually orthogonal in both the time and frequency domains, in an attempt to perturb the system in orthogonal directions.
-Similarly, pseudorandom binary sequences can be used to excite the system [11].
-The idea is to leverage the sequence’s randomness in order to excite the system in interesting ways.
-Our method differs from model-free approaches because it uses the currently estimated model to inform the next input design.
-Although they are simple, linear models are a powerful representation tool.
-They usually allow for closed-form analysis and are easily interpretable from a statistical standpoint.
-Informative input design using such models has a vast literature.
-A recent method that aims to estimate linear dynamics suggests selecting the next input by minimizing the trace of the covariance of the model parameters with linear dynamics constraints [4].
-By linearizing the objective, the authors obtain a convex problem that is easily solvable for the next input.
-This method is compatible with the dynamic mode decomposition with control (DMDc) framework that produces reduced order models for the dynamics [12].
-Our method can provide an approximate solution to any parametric model family and is exact in the case of linear models.
-In various applications, the dynamics contain modes that are not properly described by a linear model.
-In these cases, we must design inputs for nonlinear system identification.
-Early results focused on specific dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14].
-In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification.
-In this case, convex-based methods exist for informative input design [15].
-In contrast with these methods, our proposed algorithm is applicable in the case of general nonlinear models.
+This section summarizes the prior work on parameter estimation for system identification and informative input design.
+Estimating the model parameters using input-output data is usually formulated as a regression problem [9].
+A regression problem can be equivalently viewed as MLE under an appropriate distribution.
+For example, least squares regression corresponds to MLE with a Gaussian distribution.
+When estimating the model parameters, it is common to assume the true system belongs to the parametric model family [3, 4, 10].
+Then, under regularity conditions on the distribution of the input or the structure of the system, the parameter estimate is concentrated around the true system parameter [3].
+Our work deals with the problem of system identification when the true system does not necessarily belong to the selected parametric model family.
+In this case, it is critical to obtain a good fit but also a sense of the suitability of the selected model family.
+Under the assumptions mentioned in the previous paragraph, we usually obtain an unbiased estimate for the model parameters.
+This implies that we can design informative inputs by controlling the covariance of the estimator.
+Assuming an efficient estimator, the covariance converges to the inverse Fisher information matrix [4].
+Therefore, a scalar measure of the Fisher information matrix is usually optimized over the inputs.
+Different scalar functions of the Fisher information matrix are used.
+For example, a D-optimal input maximizes the determinant of the Fisher information matrix [11].
+This corresponds to minimizing the uncertainty volume of the parameter estimate.
+Our approach optimizes any desired scalar function of an approximate Fisher information matrix to determine the next input.
+Usually, the informative input design is done before the parameter estimation, but active techniques also exist [12].
+Our approach is a simple and generally applicable active system identification approach.
+In various applications, the system contains modes that are not properly described by a linear model.
+In these cases, we must deal with nonlinear models and nonlinear system identification.
+Early results focused on specific system dynamics for which the input design problem can either be relaxed using linear matrix inequalities [13] or techniques from linear systems can be applied to design informative step inputs [14].
+In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification [4].
+There is consensus among researchers that nonlinear system identification is much more difficult than linear system identification [15].
+A simple justification can be made by examining the difficulty of solving a nonlinear regression problem.
+The problem of informative input design for nonlinear dynamical systems remains a large unexplored area of research [4].
+Our approach deals with the identification of nonlinear dynamics.
+We examine methods for nonlinear system identification that are relevant to our approach.
 Goodwin presented an algorithm for informative input design in the general nonlinear case that is based on linearization of the parametric model [16].
-Our work adds to this work in an important way: we present a method to update the estimate of the model parameters that respects the linearization and is based on convex optimization, which makes it computationally fast.
-A similar work was done by Babar and Baglietto [17], who further propose an extended Kalman filter for the parameter estimation.
-We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input.
-Our objective depends on a parameter that directly encourages the satisfaction of any constraints, i.e., dynamics constraints.
-In contrast with both works [16, 17], we present a fine-tuning method for the assumed noise covariances that can improve convergence.
+Our work adds to this work in an important way: our method updates the estimate of the model parameters in a way that respects the linearization and is based on convex optimization, which makes it computationally fast.
+A similar work was done by Babar and Baglietto [17], who propose an extended Kalman filter for the joint state and parameter estimation of the dynamical system.
+Model predictive control techniques are used to select the next input.
+We improve upon their work by adding a trust-region constraint to the update of the model parameters after the linearization and by proposing a modified optimization objective for the selection of the next input that does not rely on sampling.
+In contrast with both works [16, 17], we tune the noise covariance at every iteration.
+This statistically calibrates our approach and provides information on the suitability of the selected model family for the system in question.
+We show how the approach can fail without tuning the covariance.
 III Problem Formulation
@@
 We refer to $x$ as the input and $y$ as the output.
+We treat system $g$ as a black-box.
+We can only supply an input $x$ and provide an output $y$.
 We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{q}$. At every iteration, our goal is to compute:
@@
     the optimal next input to the system that decreases our uncertainty over the unknown parameter $\theta$.
+IV Proposed Method
+We present our methodology for system identification.
 We opt for a Bayesian treatment of the problem.
-We suppose that the inputs and outputs are exactly observed, because our main goal is efficient dataset acquisition.
+We suppose that the inputs and outputs are exactly observed.
 Our method easily extends to the case of noisy observation of the outputs.
-IV Proposed Method
-We present our methodology for the design of an informative input in system identification.
-We assume that we have already observed input-output data points $(x_{i},y_{i})_{i=1}^{n}$ from the dynamics:
+We assume that we have already observed input-output data points $\{(x_{i},y_{i})\}_{i=1}^{n}$ from the system:
 $y_{i}=g(x_{i}).$ (2)
-We first outline our Bayesian approach for improving our current estimate of the model parameters.
+We first outline our Bayesian approach for improving the current estimate of the model parameters.
 We then use this Bayesian formulation in order to define an optimization problem for the next input.
-Finally, we discuss some fine-tuning that can improve performance before bringing the concepts together in our proposed algorithm.
+The components presented so far can be found in the literature [16], with the exception of the trust region in the estimate improvement step and the ability to choose the noise probability density function.
+Finally, we discuss the fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
 IV-A Parameter Estimation
-The collected data satisfy eq. 2. Because the current model $f(\cdot;\theta)$ might not match the true system $g(\cdot)$ exactly, it holds that
-$y_{i}=f(x_{i};\theta)+\epsilon_{i},$ (3)
+The collected data satisfy eq. 2. Because the current model $f(\cdot;\hat{\theta})$ might not match the true system $g(\cdot)$ exactly, it holds that
+$y_{i}=f(x_{i};\hat{\theta})+\epsilon_{i},$ (3)
 for some noise realization $\epsilon_{i}$.
-We choose the underlying noise model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma_{i})$.
+We choose the underlying noise model as $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some covariance $\Sigma$.
+Tuning this covariance is critical in the performance of the algorithm.
 Assuming conditionally independent measurements, the posterior for the hidden parameter is given by
@@
 where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$.
-The noise and prior are assumed Gaussian because they lead to a convex maximum likelihood problem [18].
-The maximum a posteriori estimate for the parameter under these assumptions is given by
+The maximum a posteriori (MAP) estimate for the parameter under these assumptions is given by
 $\hat{\theta}_{+}$ $=\operatorname*{arg\;max}_{\theta}\log p(\theta\mid y_{1},\dots,y_{n})$ (5)
 $=\operatorname*{arg\;min}_{\theta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}+$
-$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma_{i}^{-1}}^{2}.$
+$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
 In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$.
-We opt for an approximate treatment of the optimization problem that lends itself to a convex solver.
+We opt for an approximate treatment of the optimization problem that lends itself to a convex solver [18].
 Suppose that the current parameter estimate is $\hat{\theta}$.
-We do not allow our next estimate to differ much relative to the current one.
-This is justified because the next input-output pair should not change our estimate drastically.
 For model parameters close to $\hat{\theta}$, the parametric family can be accurately approximated using a first order Taylor series.
@@
 $C(x;\hat{\theta})$ $=\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}.$
-Plugging (6) into (5) and imposing the distance constraint, we get that
-$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta\lVert\hat{\theta}\rVert}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
-$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x;\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma_{i}^{-1}}^{2}.$
+Plugging (6) into (5) and imposing a trust region constraint, we get that
+$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
+$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
 This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19].
-The constraint guarantees that we search for the next model parameter in the region where the current Taylor series is accurate.
-It corresponds to a trust region.
+The constraint $\lVert\theta-\hat{\theta}\rVert\leq\delta$ guarantees that we search for the next model parameter in the region where the Taylor series for $f(\cdot;\theta)$ is sufficiently close to $f(\cdot;\theta)$.
+It also does not allow our next estimate to differ much relative to the current one.
+This is justified because the next input-output pair should not change our estimate drastically.
+Placing too much trust on any one datapoint can lead to excessively large steps or premature convergence [20].
+We can replace Gaussian densities with any other log concave density and still obtain a convex problem.
+The used density for $\epsilon_{i}$ can be changed in each estimation step, based on the realized errors $\{y_{i}-f(x_{i};\hat{\theta})\}_{i=1}^{n}$.
+This allows us to select among a family of parameter estimates.
+More details on this will be given in section IV-C.
 IV-B Informative Input Design
@@
 Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under completely linear Gaussian assumptions, the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
-$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
-It has been shown that such an expression can be accurate in the nonlinear case, although it is exact only in the linear regime [20].
+$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
+It has been shown that such an expression for the covariance can be accurate in the nonlinear case, although it is exact only in the linear regime [21].
+Eq.
+(9) is an approximate Fisher information matrix, where a prior is added and the expectation over the data is not used.
+We note that it only depends on the inputs $x_{1},\dots,x_{n+1}$ and not on the corresponding outputs.
 To minimize the posterior covariance, we maximize some notion of magnitude,
-$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
+$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
 of the posterior inverse covariance over $x_{n+1}$.
@@
 We can take such constraints into account by solving the following optimization problem:
-$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma_{i}^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
+$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
 $+\lambda\left(\mathbbm{1}(x_{n+1}\in\mathcal{X})+\mathbbm{1}(f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y})\right).$
@@
 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$.
-Equation 9 defines the posterior covariance for the parameter $\theta$. At the next iteration , we can replace the constraint of eq. 8 with
-$\lVert\theta-\hat{\theta}\rVert^{2}_{\Sigma_{\mathrm{posterior}}^{-1}}\leq\chi_{d,p}^{2},$
-where $\chi_{d,p}^{2}$ is the value of the chi-squared distribution of $d$ degrees of freedom at the $p$-percentile.
-This would allow $\theta$ to vary within the contours that correspond up to the $p$-th percentile of the posterior Gaussian distribution.
-Choosing the adequate constraint in eq.
-8 involves balancing this statistical interpretation with the accuracy of the Taylor series.
-IV-C Updating the Covariances $\Sigma_{i}$
+IV-C Finetuning
+We present the finetuning approach for the covariance $\Sigma$ that is critical for the convergence of our algorithm. We also propose a way to tune $\delta$.
 At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$.
@@
 $\hat{\epsilon}_{i}=y_{i}-f(x_{i};\hat{\theta}_{+}).$ (12)
-Using these samples, we can compute a sample covariance for the overall error in our model
-$\hat{\Sigma}_{+}=w\Sigma_{\mathrm{meas,prior}}+\dfrac{1-w}{n}\sum_{t=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top},\ w\in\left[0,1\right],$ (13)
-i.e., a convex combination of the empirical covariance and a prior covariance. This can be utilized in later iterations.
+Using these samples, we can compute a covariance for the errors caused by the mismatch between $g(\cdot)$ and $f(\cdot;\hat{\theta}_{+})$:
+$\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}.$ (13)
+However, in problem (8) we consider the distribution of the errors
+$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\theta$ (14)
+that consist of two terms: the error from the mismatch between $f(\cdot;\theta)$ and $g(\cdot)$ and the error because of the linearization, since
+$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\hat{\theta}_{+}=$ (15)
+$\hat{\epsilon}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\bar{\epsilon}_{i}}.$
+This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\}_{i=1}^{n}$:
+$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)^{\top},$ (16)
+where $\bar{\epsilon}_{i}$ is the error between the value from the model and the value from the Taylor expansion for the $i$-th output.
+For setting the trust region $\delta$, we can use a simple heuristic.
+If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$, we reduce $\delta$.
+This procedure guarantees that we keep any error due to linearization small compared to the errors because of the model mismatch.
+This allows for confident steps in the parameter update.
 IV-D Algorithm
 We combine the modules presented above in algorithm 1. We assume we have already acquired dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ and show the proposed procedure for augmenting the dataset by one point.
-Notice that when computing $x_{n+1}$ in eq.
-11 we choose $\Sigma_{i}\leftarrow\hat{\Sigma}_{+}$ from (13) for $i=1,\dots,n+1$.
-When performing the maximum a posteriori estimation in eq.
-8 , however, we use $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$.
-This selection is made because our next input should excite the system in a way that is informative and decreases the uncertainty present in our current system estimate.
-However, when improving the estimate of the model parameters we should use less informative covariances to robustify the estimation.
-Algorithm 1 Proposed Algorithm for Parameter Estimation and Input Design
-1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma_{\mathrm{meas,prior}}$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta,\lambda$ and $w$
-2:$\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}},\ \mathrm{for\ all}\ i=1,\dots,n$
+Our algorithm first runs a loop to improve the parameter estimate.
+We obtain a new parameter estimate by solving problem (8).
+We only accept the new solution if the linearization errors are insignificant compared to the model errors.
+In this case, we update the covariance $\Sigma$ and resolve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate.
+If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and resolve.
+This iterative technique will produce a new parameter estimate.
+It will further produce an accurate error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
+In the second phase, we obtain the next input by optimizing eq. 11. Note that we use $\Sigma_{\mathrm{model\ error}}$, because we aim to reduce the uncertainty associated with the model mismatch.
+The algorithm outputs an updated $\Sigma$, which is used as input in the next algorithm call.
+Algorithm 1 allows for testing whether the chosen parametric model family $f$ is suitable for the system under consideration.
+We know that the parameter estimates obtained from the successive calls to algorithm 1 correspond to the best model fits for the obtained datapoints and Taylor MAP objective.
+Suppose that the magnitude of $\Sigma_{\mathrm{model\ error}}$ does not decrease between successive calls to algorithm 1.
+This suggests that the chosen parametric family $f$ is not appropriate for the system.
+Algorithm 1 Proposed Algorithm for System Identification
+1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta$ and $\lambda$
+2:repeat
 3:     Obtain updated estimate $\hat{\theta}_{+}$ by eq. 8
-4:$\Sigma_{i}\leftarrow\hat{\Sigma}_{+},\ \mathrm{for\ all}\ i=1,\dots,n+1$
-5:Obtain next input $x_{n+1}$ by solving eq. 11
-6:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
-7:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$
+4:     Compute $\hat{\epsilon}_{i},\bar{\epsilon}_{i}$ for all $i=1,\dots,n$
+5:     if $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$ then
+6:         Reduce $\delta$
+7:     else
+8:         Update $\Sigma$ using eq. 16
+9:         Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}$
+10:         $\hat{\theta}\leftarrow\hat{\theta}_{+}$
+11:until convergence
+12:Obtain next input $x_{n+1}$ by solving eq. 11 where $\Sigma\leftarrow\Sigma_{\mathrm{model\ error}}$
+13:Obtain output $y_{n+1}$ for input $x_{n+1}$ using the true system
+14:return dataset $\{(x_{i},y_{i})\}_{i=1}^{n+1}$, $\hat{\theta}_{+}$, $\Sigma$, and $\Sigma_{\mathrm{model\ error}}$
 V Results
 We test our algorithm on both linear and nonlinear systems.
-[^1] The code is included in the repository: https://github.com/alextzik/informative_input_design We further perform an ablation study to show the importance of finetuning the covariances $\Sigma_{i}$ in our algorithm.
+[^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of finetuning the covariances $\Sigma$ in our algorithm.
 Baselines
-Our emphasis is on the informative input design.
-We compare against three different model-free baselines for the selection of inputs.
-We maintain the same parameter estimation module from section IV in all methods.
-The first baseline is random.
-At every iteration, we run algorithm 1, with a change in line 4: we suppose $x_{n+1}\sim\mathcal{N}(0,I)$, i.e., $x_{n+1}$ is randomly selected.
-The second baseline is based on pseudorandom binary sequences and is called PRBS.
-At every iteration, we run algorithm 1, with a change in line 4: we suppose that each coordinate of $x_{n+1}$ originates from a pseudorandom binary sequence.
-Finally, the last baseline is called mutlisine and is based on orthogonal multisine inputs [10].
-In line 4 of algorithm 1, we set the coordinate $i$ of $x_{n+1}$ as follows
-$x_{n+1,i}=\sum_{j=0}^{M-1}A_{i,j}\sin\left(\dfrac{2\pi(d_{x}j+i)(n+1)}{T}+\phi_{i,j}\right),$
-where $M$ is the total number of sinusoidal components and $T$ is the total number of inputs to be designed.
-We sample each phase $\phi_{i,j}$ uniformly in $\left[0,2\pi\right]$ and each amplitude $A_{i,j}$ uniformly in $\left[-1,1\right]$.
-We select $M=30$ in all experiments below.
-Algorithm Parameters
-We set $w=0.2$, $\delta=0.3$, and $M=30$.
+We consider a variant of our proposed algorithm as baseline.
+The baseline does not rely on any finetuning.
+Given the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, $\Sigma$, $\Sigma_{\mathrm{prior}}$ and $\hat{\theta}$, it obtains an updated estimate $\hat{\theta}_{+}$ by solving eq.
+8.
+It then solves eq.
+11 to obtain the next input $x_{n+1}$.
+It does not update $\Sigma$, which is always equal to the initial value provided.
+Set-Up
+We set $\delta=0.3$.
+In line 5 of algorithm 1, we reduce $\delta$ by setting $\delta\leftarrow 0.8\delta$.
 We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition.
-We do set an upper bound on the allowed input magnitude.
-If the input provided by any method has a larger magnitude, it is scaled to have the largest allowed magnitude.
-The first datapoint is assumed fixed for all methods.
+The first datapoints are assumed fixed for all methods.
+Further, we run the algorithms with multiple initializations of $\Sigma$.
+We use the scalar function $\log\det$ as $\mathcal{M}$ when determining the next input.
 Metrics
-In all cases below , the true system belongs to the parametric model family.
+In most cases below , the true system belongs to the parametric model family.
 We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number.
@@
 9.
-We assume $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$.
 A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs.
@@
 Our results are shown in Figure 3.
+In this experiment, we also bound the norm of the input to be less than $0.5$.
 In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses.
 We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional.
-We further plot the designed inputs for the various methods.
-Our algorithm exhibits a faster convergence to the true model parameters compared to the baselines.
+We further plot the designed inputs for the two methods.
+Our algorithm exhibits a faster convergence to the true model parameters compared to the baseline.
 This is verified by the lower value of $\log\det$ for the final error covariance.
@@
 Our results are shown in Figure 4.
-Our proposed algorithm is able to converge to the true model parameters and outperforms the baselines with respect to both metrics considered.
-From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter the estimate has a higher uncertainty in the second coordinate $\theta_{2}$, which is captured by the elongated ellipse axes along that direction.
+Our proposed algorithm is able to converge to the true model parameters and outperforms the baseline with respect to both metrics considered.
+From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter value, the estimate along its trajectory has a higher uncertainty in the first coordinate $\theta_{1}$, which is captured by the elongated ellipse axis along that direction.
+We also observe that the error ellipses shrink as we move towards the true parameters.
+This implies that our uncertainty is reduced.
 V-C Unicycle Model
@@
 The time discretization is thus one of the unknown parameters.
-We start the model from the origin with zero being the first control for all methods.
+We start the system from the origin with zero being the first control for both methods.
 Figure 5 shows the performance of our algorithm.
-It outperforms the considered baselines.
-We further plot $(x_{t,1},x_{t,2})$ for the various methods.
-V-D Ablation
-In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with a variant that uses $\Sigma_{i}\leftarrow\Sigma_{\mathrm{meas,prior}}$ throughout.
-The new variant is termed proposed (no finetuning).
-In this simulation , we remove the amplitude constraint.
-It is evident in Figure 1 that the variant does not converge properly.
-This highlights the importance of finetuning the covariances $\Sigma_{i}$.
-Figure 1: Ablation study in the case of Hénon dynamics.
-(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
-(b) Model parameter estimate error over timesteps.
-(c) Inputs as a function of timestep for the various methods.
+It outperforms the considered baseline.
+The posterior error ellipses are an accurate representation of the uncertainty.
+Once, $\theta_{1}$ is correctly estimated, the ellipses become vertical, since there is uncertainty only in $\theta_{2}$.
+We further plot the state trajectory $(x_{t,1},x_{t,2})$ for the two methods.
+In this experiment, we bound the norm of the input to be less than $1$.
+V-D Importance of Finetuning
+In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with the baseline.
+The initial covariance $\Sigma$ is set to $0.2I$, while $\Sigma_{\mathrm{prior}}=0.1I$ and $\theta_{\mathrm{prior}}=(7,4)$.
+It is evident in Figure 1 that the baseline does not converge properly.
+This can be attributed to the high confidence we have initially placed on both the prior value of $\theta$ and the accuracy of the measurements.
+Under these conditions, without finetuning, it is impossible for the algorithm to detect the false confidence.
+This highlights the importance of finetuning the covariance $\Sigma$.
+Figure 1: Importance of finetuning the covariance matrix $\Sigma$ in the case of Hénon dynamics.
+V-E Model Mismatch
+We consider the case where the parametric model family does not include the true system.
+We suppose that the true system is the Hénon dynamics of eq.
+19 and the parametric model family consists of linear models of the form
+$f(x_{t};\theta)=\begin{bmatrix}\theta_{1}&\theta_{2}\\ \theta_{1}&\theta_{2}\end{bmatrix}x_{t}.$ (23)
+Obviously, this is an inadequate model family.
+As discussed in section IV-D, we expect our algorithm to provide an indication of the inadequacy of the parametric model family via the magnitude of $\Sigma_{\mathrm{model\ error}}$.
+This is shown in Figure 2.
+The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease as the iterations progress.
+This implies that the realized errors remain high.
+Given that our algorithm computes MLEs, this indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics.
+Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over iterations when the parametric model family cannot capture the true system dynamics.
+The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease indicates an inadequate parametric model family.
+(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
+(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
 Figure 3: Results for the linear dynamics.
-(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
-(b) Model parameter estimate error over timesteps.
-(c) Inputs as a function of timestep for the various methods.
+(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
+(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
 Figure 4: Results for the Hénon dynamics.
-(a) Model parameter estimate in proposed algorithm as it evolves over timesteps.
-(b) Model parameter estimate error over timesteps.
-(c) State of dynamics as a function of timestep for the various methods.
+(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
+(b) State of dynamics as a function of time for the various methods for a specific initialization of $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
 Figure 5: Results for the unicycle dynamics.
 Method/Experiment Linear Dynamics Hénon Dynamics Unicycle Dynamics
-proposed $\mathbf{-18.5}$ $\mathbf{-14.9}$ $\mathbf{-9.8}$
-random $-10.7$ $-5.9$ $-7.0$
-PRBS $-10.2$ $-5.1$ $-9.2$
-multisine $-13.7$ $-9.7$ $-8.9$
+proposed $\mathbf{-77.2}$ $\mathbf{-51.2}$ $\mathbf{-43.5}$
+baseline $-12.1$ $-31.8$ $-7.5$
 TABLE I: $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep for the various dynamics studied.
 VI Conclusion
-We present a Bayesian methodology for informative input design that is based on linear Gaussian assumptions.
+We present a Bayesian methodology for system identification that is based on linear Gaussian assumptions.
 It extends to general nonlinear system identification through the Taylor expansion.
-A key module in our algorithm is the online calibration of the noise covariances that greatly impacts performance.
+A key module in our algorithm is the online calibration of the noise covariance that greatly impacts performance and allows us to judge the suitability of a parametric model family for the underlying system.
 Our algorithm also has limitations.
-We must select the trust region shape and volume when performing the parameter estimation , and we do not address the problem of selecting the parametric model family.
-Future work will propose a method for judging the suitability of a given parametric model family.
+The use of the trust region, although necessary to ensure the accuracy of the Taylor series, might forbid large changes in the parameters when these are implied by the data.
+Future work will apply our method to high-dimensional problems and look into theoretical guarantees.
 Acknowledgements

### v3

@@
 We tackle the problem of system identification, where we select inputs, observe the corresponding outputs from the true system, and optimize the parameters of our model to best fit the data.
-We propose a flexible and computationally tractable methodology that is compatible with any system and parametric family of models.
-Our approach only requires input-output data from the system and first-order information from the model with respect to the parameters.
-Our algorithm consists of two modules.
+We propose a practical and computationally tractable methodology that is compatible with any system and parametric family of models.
+Our approach only requires input-output data from the system and first-order information of the model with respect to the parameters.
+Our approach consists of two modules.
 First, we formulate the problem of system identification from a Bayesian perspective and use a linear Gaussian model approximation to iteratively optimize the model’s parameters.
 In each iteration, we propose to use the input-output data to tune the covariance of the linear Gaussian model.
-This statistically calibrates the approach.
+This online covariance calibration stabilizes fitting and signals model inaccuracy.
 Secondly, we define a Gaussian-based uncertainty measure for the model parameters, which we can then minimize with respect to the next selected input.
@@
 Fitting model parameters to data can be challenging when the model is a complicated function of the parameters.
-In addition, Ott et al.
+Ott et al.
 have shown that selecting control inputs based on the current model can improve the system identification process [2].
-Informative input design can reduce the data gathering requirements for the purpose of modeling the system, which can be costly in real-world applications.
+Informative input design can reduce the data gathering requirements for system modeling, which can be costly in real-world applications.
 This can be particularly useful in high-dimensional problems, where finding informative directions to perturb the system can be challenging.
@@
 The family must be expressive enough to capture the true system, but also simple enough to optimize over.
-Our parameter estimate is improved once more data becomes available and we must also choose the optimization approach that updates the model parameters.
+Our parameter estimate is improved once more data becomes available.
+We must also choose the optimization approach that updates the model parameters.
 A vast literature exists on methods for estimating the model parameters [3].
@@
 Our work deals with the following question:
-  • Can we design a method for system identification that is computationally tractable, compatible with any parametric model family and able to suggest whether a parametric model family is suitable for a given system?
+  • Can we design a method for system identification that is computationally light, compatible with any parametric model family and able to determine the suitability of the model family for a given system?
+Such a method would be useful for practitioners who usually want a fast and model-agnostic approach.
 Our approach leverages first-order information to construct a linear Gaussian model for the output given the input, as a function of the model parameters.
-Based on the first-order Taylor expansion, we solve successive convex problems, corresponding to maximum likelihood estimation (MLE), to update the model parameters.
+We use the first-order Taylor expansion with respect to the model parameters and solve successive convex problems to improve the model parameter estimate.
 This simplifies the model update step discussed in the previous paragraph, which becomes computationally tractable.
-We also propose a procedure to tune the covariance associated with the linear Gaussian model.
+We also propose a procedure to tune the covariance associated with the linear Gaussian model in an online manner.
 This allows us to statistically calibrate our model and judge the suitability of the parametric model family.
-Finally, using the linear Gaussian model from the first-order Taylor expansion, we define an uncertainty measure that is accurate in the linear Gaussian case for the input design.
+Finally, using the linear Gaussian model we define an uncertainty measure to optimize over the next input.
 The rest of the paper is organized as follows.
@@
 The problem of system identification has been well studied.
-Early research on the optimal selection of inputs dates back to the 1960s [5, 6] and highlights the importance of the problem for the identification of systems.
+Research on the optimal selection of inputs dates back to the 1960s [5, 6].
 Recent advances in computational capabilities and learning-based methods have once again brought attention to the problem [7, 8].
-This section summarizes the prior work on parameter estimation for system identification and informative input design.
+This section summarizes the prior work on parameter estimation and informative input design for system identification.
 Estimating the model parameters using input-output data is usually formulated as a regression problem [9].
-A regression problem can be equivalently viewed as MLE under an appropriate distribution.
+A regression problem can be equivalently viewed as maximum likelihood estimation (MLE) under an appropriate distribution.
 For example, least squares regression corresponds to MLE with a Gaussian distribution.
@@
 In certain instances, the nonlinearity can be separated from the linear part of the model in what is known as structured nonlinear system identification [4].
-There is consensus among researchers that nonlinear system identification is much more difficult than linear system identification [15].
-A simple justification can be made by examining the difficulty of solving a nonlinear regression problem.
-The problem of informative input design for nonlinear dynamical systems remains a large unexplored area of research [4].
-Our approach deals with the identification of nonlinear dynamics.
+There is consensus among researchers that nonlinear system identification is more difficult than linear system identification [15].
+This is, to some extent, because nonlinear optimization is more challenging than linear programming.
+Our approach deals with the identification of nonlinear dynamics using convex programming.
+Convex programs are solved efficiently using existing packages [16].
 We examine methods for nonlinear system identification that are relevant to our approach.
@@
 In contrast with both works [17, 18], we tune the noise covariance at every iteration.
-This statistically calibrates our approach and provides information on the suitability of the selected model family for the system in question.
-We show how the approach can fail without tuning the covariance.
+In addition to statistically calibrating our approach, the tuned covariance serves as a lightweight diagnostic about the suitability of the selected model family.
+We demonstrate this empirically.
 III Problem Formulation
@@
 where $x\in\mathbb{R}^{d_{x}},y\in\mathbb{R}^{d_{y}}$ and $g:\mathbb{R}^{d_{x}}\rightarrow\mathbb{R}^{d_{y}}$.
+We refer to $x$ as the input and $y$ as the output.
 If $g$ represents dynamics, the input is usually the current state and control, which we abstract as $x$, while the next state is abstracted as $y$.
-We refer to $x$ as the input and $y$ as the output.
 We treat system $g$ as a black-box.
-We can only supply an input $x$ and provide an output $y$.
-We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{q}$. At every iteration, our goal is to compute:
+We can only supply an input $x$ and obtain an output $y$ from the system.
+We choose a family of functions $f(\cdot;\theta)$, parameterized by $\theta\in\mathbb{R}^{d_{\theta}}$. At every iteration, our goal is to compute:
   1. 1.
-    $\hat{\theta}$ such that $g(\cdot)\approx f(\cdot;\hat{\theta})$ using the available data,
+[…]
   2. 2.
     the optimal next input to the system that decreases our uncertainty over the unknown parameter $\theta$.
+In the present work, we do not deal with the selection process for the parametric family $f(\cdot;\theta)$.
+Nevertheless, as we will see below, our approach allows the user to test its suitability for the underlying system.
 IV Proposed Method
@@
 We then use this Bayesian formulation in order to define an optimization problem for the next input.
-The components presented so far can be found in the literature [16], with the exception of the trust region in the estimate improvement step and the ability to choose the noise probability density function.
-Finally, we discuss the fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
+The components presented so far can be found in the literature [17], with the exception of the trust region in the estimate improvement step and the ability to choose the underlying probability density functions.
+Finally, we discuss the covariance fine-tuning, which is critical in the convergence of the method and constitutes our major contribution.
 IV-A Parameter Estimation
-The collected data satisfy eq. 2. Because the current model $f(\cdot;\hat{\theta})$ might not match the true system $g(\cdot)$ exactly, it holds that
+[…]
 $y_{i}=f(x_{i};\hat{\theta})+\epsilon_{i},$ (3)
-for some noise realization $\epsilon_{i}$.
-We choose the underlying noise model as $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some covariance $\Sigma$.
+where $\epsilon_{i}$ is the model error for the $i$-th data-point.
+We treat $\epsilon_{i}$ as a random quantity and choose the underlying model $\epsilon_{i}\sim\mathcal{N}(0,\Sigma)$ for some positive definite covariance $\Sigma$.
 Tuning this covariance is critical in the performance of the algorithm.
@@
 $p(\theta\mid y_{1},\dots,y_{n})\propto p(\theta)\prod_{i=1}^{n}p(y_{i}\mid\theta),$ (4)
-where we choose the prior $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$.
-The maximum a posteriori (MAP) estimate for the parameter under these assumptions is given by
+where we choose the prior for $\theta$, $p(\theta)$.
+The maximum a-posteriori (MAP) estimate for the parameter under these assumptions is given by
 $\hat{\theta}_{+}$ $=\operatorname*{arg\;max}_{\theta}\log p(\theta\mid y_{1},\dots,y_{n})$ (5)
-$=\operatorname*{arg\;min}_{\theta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}+$
-$\qquad\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
+$=\operatorname*{arg\;min}_{\theta}\ -\log p(\theta)+\dfrac{1}{2}\sum_{i=1}^{n}\lVert y_{i}-f(x_{i};\theta)\rVert_{\Sigma^{-1}}^{2}.$
 In general, this can be a very difficult problem to solve because it depends on the structure of the parametric family $f$ with respect to the parameter $\theta$.
+We can look for a local optimum using gradient-based techniques.
 We opt for an approximate treatment of the optimization problem that lends itself to a convex solver [19].
@@
 We write
-$f(x;\theta)$ $\approx f(x;\hat{\theta})+\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\left(\theta-\hat{\theta}\right)$ (6)
+[…]
 $=b(x;\hat{\theta})+C(x;\hat{\theta})\theta$
@@
 $\left[\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\right]_{i,j}$ $=\dfrac{\partial f_{i}}{\partial\theta_{j}}|_{(x;\hat{\theta})}$, (7)
-$b(x;\hat{\theta})$ $=f(x;\hat{\theta})-\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}\hat{\theta},$
-$C(x;\hat{\theta})$ $=\dfrac{\partial f}{\partial\theta}|_{(x;\hat{\theta})}.$
+[…]
+[…]
 Plugging (6) into (5) and imposing a trust region constraint, we get that
-$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}\ \dfrac{1}{2}\lVert\theta-\theta_{\mathrm{prior}}\rVert^{2}_{\Sigma_{\mathrm{prior}}^{-1}}$ (8)
-$+\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
-This is a convex problem with respect to $\theta$ and can easily be handled by a computational framework for convex optimization [19].
-The constraint $\lVert\theta-\hat{\theta}\rVert\leq\delta$ guarantees that we search for the next model parameter in the region where the Taylor series for $f(\cdot;\theta)$ is sufficiently close to $f(\cdot;\theta)$.
-It also does not allow our next estimate to differ much relative to the current one.
-This is justified because the next input-output pair should not change our estimate drastically.
-Placing too much trust on any one datapoint can lead to excessively large steps or premature convergence [20].
+$\hat{\theta}_{+}\approx\operatorname*{arg\;min}_{\lVert\theta-\hat{\theta}\rVert\leq\delta}$ $-\log p(\theta)+$ (8)
+$\dfrac{1}{2}\sum_{i=1}^{n}\lVert\underbrace{y_{i}-b(x_{i};\hat{\theta})}_{\tilde{y}_{i}}-C(x_{i};\hat{\theta})\theta\rVert_{\Sigma^{-1}}^{2}.$
+[…]
 We can replace Gaussian densities with any other log concave density and still obtain a convex problem.
 The used density for $\epsilon_{i}$ can be changed in each estimation step, based on the realized errors $\{y_{i}-f(x_{i};\hat{\theta})\}_{i=1}^{n}$.
-This allows us to select among a family of parameter estimates.
+This allows us to select the family of underlying distributions.
 More details on this will be given in section IV-C.
@@
 The goal of selecting $x_{n+1}$ is to reduce our uncertainty in the posterior estimate of $\theta$.
-Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under completely linear Gaussian assumptions, the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
-$\Sigma_{\mathrm{posterior}}^{-1}=\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+}).$ (9)
-It has been shown that such an expression for the covariance can be accurate in the nonlinear case, although it is exact only in the linear regime [21].
+Assuming that we are given a new data-point $(x_{n+1},y_{n+1})$ and under linear Gaussian assumptions (for the prior as well), the posterior inverse covariance is given by the Hessian of the objective of (8) but evaluated using the most up-to-date MAP estimate $\hat{\theta}_{+}$, i.e,
+[…]
+where we assume $\theta\sim\mathcal{N}(\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}})$.
+In the case of a uniform prior for $\theta$, we can set $\Sigma_{\mathrm{prior}}^{-1}=0$, i.e., we can simply ignore the prior term in (9).
+It has been shown that the covariance (9) can be accurate in the nonlinear case, although it is exact only in the linear regime [21].
 Eq.
@@
 To minimize the posterior covariance, we maximize some notion of magnitude,
-$\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right),$ (10)
+[…]
 of the posterior inverse covariance over $x_{n+1}$.
@@
 We can take such constraints into account by solving the following optimization problem:
-$\operatorname*{arg\;max}_{x_{n+1}}$ $\mathcal{M}\left(\Sigma_{\mathrm{prior}}^{-1}+\sum_{i=1}^{n+1}C(x_{i};\hat{\theta}_{+})^{\top}\Sigma^{-1}C(x_{i};\hat{\theta}_{+})\right)$ (11)
+[…]
 $+\lambda\left(\mathbbm{1}(x_{n+1}\in\mathcal{X})+\mathbbm{1}(f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y})\right).$
@@
 11 to optimize over the next $k$ inputs, $x_{n+1},\dots,x_{n+k}$.
-IV-C Finetuning
-We present the finetuning approach for the covariance $\Sigma$ that is critical for the convergence of our algorithm. We also propose a way to tune $\delta$.
+In (11), because the system $g$ is only accessible via a query oracle, we approximate $g(x_{n+1})\in\mathcal{Y}$ by $f(x_{n+1};\hat{\theta}_{+})\in\mathcal{Y}$.
+IV-C Fine-tuning
+We present the fine-tuning approach for the covariance $\Sigma$ and also propose a way to tune $\delta$.
 At the current iteration , we have access to the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$.
-After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\hat{\epsilon}_{i}$ by comparing the true output to the expected output, i.e., for $i=1,\dots,n$
-$\hat{\epsilon}_{i}=y_{i}-f(x_{i};\hat{\theta}_{+}).$ (12)
-Using these samples, we can compute a covariance for the errors caused by the mismatch between $g(\cdot)$ and $f(\cdot;\hat{\theta}_{+})$:
-$\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}.$ (13)
+After obtaining the new parameter estimate $\hat{\theta}_{+}$, we can compute the estimated noise realizations $\epsilon^{\mathrm{model}}_{i}$ by comparing the true output to the model output, i.e., for $i=1,\dots,n$
+[…]
+[…]
+$\Sigma_{\mathrm{model\ error}}=:\dfrac{1}{n}\sum_{i=1}^{n}\epsilon^{\mathrm{model}}_{i}{\epsilon^{\mathrm{model}}_{i}}^{\top}.$ (13)
 However, in problem (8) we consider the distribution of the errors
-$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\theta$ (14)
+$y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\hat{\theta}_{+}$ (14)
 that consist of two terms: the error from the mismatch between $f(\cdot;\theta)$ and $g(\cdot)$ and the error because of the linearization, since
 $y_{i}-b(x_{i};\hat{\theta})-C(x_{i};\hat{\theta})\hat{\theta}_{+}$= (15)
-$\hat{\epsilon}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\bar{\epsilon}_{i}}.$
-This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\}_{i=1}^{n}$:
-$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)\left(\hat{\epsilon}_{i}+\bar{\epsilon}_{i}\right)^{\top},$ (16)
-where $\bar{\epsilon}_{i}$ is the error between the value from the model and the value from the Taylor expansion for the $i$-th output.
+$\epsilon^{\mathrm{model}}_{i}+\underbrace{f(x_{i};\hat{\theta}_{+})-\left(b(x_{i};\hat{\theta})+C(x_{i};\hat{\theta})\hat{\theta}_{+}\right)}_{\epsilon^{\mathrm{lin}}_{i}}.$
+This implies that the covariance $\Sigma$ should be set using the sample covariance of $\{\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\}_{i=1}^{n}$:
+$\Sigma\leftarrow\dfrac{1}{n}\sum_{i=1}^{n}\left(\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\right)\left(\epsilon^{\mathrm{model}}_{i}+\epsilon^{\mathrm{lin}}_{i}\right)^{\top},$ (16)
+where $\epsilon^{\mathrm{lin}}_{i}$ is the error between the value from the model and the value from the Taylor expansion of the model for the $i$-th output.
 For setting the trust region $\delta$, we can use a simple heuristic.
-If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$, we reduce $\delta$.
+If the average linearization error size $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{lin}}_{i}\rVert$ becomes comparable to the average absolute model error size $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{model}}_{i}\rVert$, we reduce $\delta$.
 This procedure guarantees that we keep any error due to linearization small compared to the errors because of the model mismatch.
@@
 We obtain a new parameter estimate by solving problem (8).
-We only accept the new solution if the linearization errors are insignificant compared to the model errors.
-In this case, we update the covariance $\Sigma$ and resolve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate.
-If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and resolve.
+We only accept the new solution if the linearization errors are small compared to the model errors.
+In this case, we update the covariance $\Sigma$ and re-solve problem (8) using the new covariance and the Taylor expansion around the new parameter estimate.
+If the linearization errors are deemed significant we reject the new estimate, reduce $\delta$ and re-solve.
 This iterative technique will produce a new parameter estimate.
-It will further produce an accurate error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
+It will further produce an error covariance $\Sigma_{\mathrm{model\ error}}$ for the errors attributed to the discrepancy between the current model and the true system.
 In the second phase, we obtain the next input by optimizing eq. 11. Note that we use $\Sigma_{\mathrm{model\ error}}$, because we aim to reduce the uncertainty associated with the model mismatch.
@@
 Algorithm 1 allows for testing whether the chosen parametric model family $f$ is suitable for the system under consideration.
-We know that the parameter estimates obtained from the successive calls to algorithm 1 correspond to the best model fits for the obtained datapoints and Taylor MAP objective.
+We know that the parameter estimates obtained from the successive calls to Algorithm 1 correspond to the best model fits for the obtained data-points and the linearized MAP objective.
 Suppose that the magnitude of $\Sigma_{\mathrm{model\ error}}$ does not decrease between successive calls to Algorithm 1.
 This suggests that the chosen parametric family $f$ is not appropriate for the system.
-Algorithm 1 Proposed Algorithm for System Identification
+Algorithm 1 Proposed Algorithm for Active System Identification
 1:dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, parameter estimate $\hat{\theta}$, covariances $\Sigma_{\mathrm{prior}}$ and $\Sigma$, model family $f$, constraint sets $\mathcal{X},\mathcal{Y}$, constants $\delta$ and $\lambda$
@@
 3:  Obtain updated estimate $\hat{\theta}_{+}$ by eq. 8
-4:     Compute $\hat{\epsilon}_{i},\bar{\epsilon}_{i}$ for all $i=1,\dots,n$
-5:     if $n^{-1}\sum_{i=1}^{n}\lVert\bar{\epsilon}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\hat{\epsilon}_{i}\rVert$ then
+4:  Compute $\epsilon^{\mathrm{model}}_{i},\epsilon^{\mathrm{lin}}_{i}$ for all $i=1,\dots,n$
+5:  if $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{lin}}_{i}\rVert$ large compared to $n^{-1}\sum_{i=1}^{n}\lVert\epsilon^{\mathrm{model}}_{i}\rVert$ then
 6:   Reduce $\delta$
@@
 8:   Update $\Sigma$ using eq. 16
-9:         Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\hat{\epsilon}_{i}\hat{\epsilon}_{i}^{\top}$
+9:   Set $\Sigma_{\mathrm{model\ error}}=\dfrac{1}{n}\sum_{i=1}^{n}\epsilon^{\mathrm{model}}_{i}{\epsilon^{\mathrm{model}}_{i}}^{\top}$
 10:   $\hat{\theta}\leftarrow\hat{\theta}_{+}$
@@
 We test our algorithm on both linear and nonlinear systems.
-[^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of finetuning the covariances $\Sigma$ in our algorithm.
-Baselines
-We consider a variant of our proposed algorithm as baseline.
-The baseline does not rely on any finetuning.
-Given the dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$, $\Sigma$, $\Sigma_{\mathrm{prior}}$ and $\hat{\theta}$, it obtains an updated estimate $\hat{\theta}_{+}$ by solving eq.
-8.
-It then solves eq.
-11 to obtain the next input $x_{n+1}$.
-It does not update $\Sigma$, which is always equal to the initial value provided.
+[^1] The code is included in the repository: https://github.com/alextzik/informative_input_design Our results show the importance of fine-tuning the covariances $\Sigma$ in our algorithm.
 Set-up
@@
 In line 5 of algorithm 1, we reduce $\delta$ by setting $\delta\leftarrow 0.8\delta$.
-We do not consider any constraint s on $\mathcal{X}$ or $\mathcal{Y}$ for simplicity in the exposition.
-The first datapoints are assumed fixed for all methods.
-Further, we run the algorithms with multiple initializations of $\Sigma$.
+The first data-points are assumed fixed.
+Furthermore, we independently run the algorithm for $30$ different initializations of $\Sigma$, $\theta_{\mathrm{prior}}$, and $\Sigma_{\mathrm{prior}}$.
 We use the scalar function $\log\det$ as $\mathcal{M}$ when determining the next input.
+The loop in lines 1-10 in Algorithm 1 is carried out for $10$ times.
 Metrics
-In most cases below , the true system belongs to the parametric model family.
+In section V-A, section V-B, and section V-C below , the true system belongs to the parametric model family, i.e., $g(\cdot)=f(\cdot;\theta_{\mathrm{true}})$ for some $\theta_{\mathrm{true}}\in\mathbb{R}^{d_{\theta}}$.
 We can thus compare the methods with respect to the error $\lVert\hat{\theta}-\theta_{\mathrm{true}}\rVert_{\infty}$ as a function of iteration number.
-Further, we can compare the methods with respect to the $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep, which is given in eq.
-9.
-A lower value signifies less uncertainty in the parameter estimate and a more informative sequence of inputs.
-The results are gathered in Table I for all experiments.
+The iteration equals the size of the dataset.
 V-A Linear System
@@
 In this experiment, we also bound the norm of the input to be less than $0.5$.
-In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses.
-We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, although the complete estimate is 4-dimensional.
-We further plot the designed inputs for the two methods.
-Our algorithm exhibits a faster convergence to the true model parameters compared to the baseline.
-This is verified by the lower value of $\log\det$ for the final error covariance.
+Based on the error plot, our algorithm is able to converge to the true parameter values irrespective of the initial $\Sigma$, $\theta_{\mathrm{prior}}$, and $\Sigma_{\mathrm{prior}}$.
+In addition to the error, we plot the trajectory of the parameter estimate along with its error ellipses for a particular choice of $\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}}$, and initial $\Sigma$.
+We only show the estimate d trajectory of the first two coordinates of $\hat{\theta}$, $\hat{\theta}_{1}$ and $\hat{\theta}_{2}$, although the complete estimate is 4-dimensional.
+We further plot the designed inputs for the same $\theta_{\mathrm{prior}},\Sigma_{\mathrm{prior}}$, and initial $\Sigma$.
 V-B Hénon map dynamics
@@
 Our results are shown in Figure 4.
-Our proposed algorithm is able to converge to the true model parameters and outperforms the baseline with respect to both metrics considered.
-From the estimate trajectory, we observe that the covariance truly captures the uncertainty in the parameter: based on the true parameter value, the estimate along its trajectory has a higher uncertainty in the first coordinate $\theta_{1}$, which is captured by the elongated ellipse axis along that direction.
+Our proposed algorithm is able to converge to the true model parameters.
+From the estimate trajectory, we observe that the posterior covariance (9) captures the uncertainty in the parameter: based on the true parameter value, the estimate has larger errors in the first coordinate $\theta_{1}$ at later times, which is captured by the elongated ellipse axis along that direction for those times.
 We also observe that the error ellipses shrink as we move towards the true parameters.
@@
 The parametric model is assumed to be
-$f(x_{t},u_{t};\theta)=\begin{bmatrix}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}.$ (22)
+$f(x_{t},u_{t};\theta)=\begin{bmatrix}\theta_{2}x_{t,1}+u_{1,t}\theta_{1}\cos(x_{t,3})\\ \theta_{2}x_{t,2}+u_{1,t}\theta_{1}\sin(x_{t,3})\\ x_{t,3}+u_{2,t}\theta_{1}\end{bmatrix}$. (22)
 The time discretization is thus one of the unknown parameters.
-We start the system from the origin with zero being the first control for both methods.
 Figure 5 shows the performance of our algorithm.
-It outperforms the considered baseline.
-The posterior error ellipses are an accurate representation of the uncertainty.
 Once, $\theta_{1}$ is correctly estimated, the ellipses become vertical, since there is uncertainty only in $\theta_{2}$.
-We further plot the state trajectory $(x_{t,1},x_{t,2})$ for the two methods.
-In this experiment, we bound the norm of the input to be less than $1$.
-V-D Importance of Finetuning
-In order to show the significance of finetuning, we compare, in the case of the Hénon dynamics, algorithm 1 with the baseline.
-The initial covariance $\Sigma$ is set to $0.2I$, while $\Sigma_{\mathrm{prior}}=0.1I$ and $\theta_{\mathrm{prior}}=(7,4)$.
-It is evident in Figure 1 that the baseline does not converge properly.
-This can be attributed to the high confidence we have initially placed on both the prior value of $\theta$ and the accuracy of the measurements.
-Under these conditions, without finetuning, it is impossible for the algorithm to detect the false confidence.
-This highlights the importance of finetuning the covariance $\Sigma$.
-Figure 1: Importance of finetuning the covariance matrix $\Sigma$ in the case of Hénon dynamics.
-V-E Model Mismatch
+In this experiment, we bound the norm of the input $u_{1,t}$ to be less than $1$.
+V-D Model Mismatch
 We consider the case where the parametric model family does not include the true system.
+We say that the model family does not capture the true dynamics.
 We suppose that the true system is the Hénon dynamics of eq.
@@
 This is shown in Figure 1.
-The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease as the iterations progress.
+The magnitude of $\Sigma_{\mathrm{model\ error}}$, here chosen as $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value.
 This implies that the realized errors remain high.
-Given that our algorithm computes MLEs, this indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics.
-Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over iterations when the parametric model family cannot capture the true system dynamics.
-The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ does not decrease indicates an inadequate parametric model family.
-(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
-(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
-(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
+Lines 4-5 of Algorithm 1 ensure that the large model errors are not because of a bad step in the optimization of (8).
+In other words, lines 4-5 make it unlikely that we have the appropriate model family, but a bad estimate $\hat{\theta}$ due to the linearization error in problem (8).
+This indicates that the obtained parameter estimates, which are in a sense optimal, are inadequate to capture the true dynamics.
+In the case of the model family (20), which captures the dynamics, $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus lower than its initial value.
+In other words, we suggest the following empirical guideline:
+If $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus at a value higher than its initial value and sufficiently large (close to 0) or rises across time, consider modifying the model family.
+Equivalently, we can look at the model errors $\epsilon_{i}^{\mathrm{model}}$ to draw our conclusions.
+We repeat the process for the inadequate model family (18) in Figure 2.
+Figure 1: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over time for the parametric model family (23) and the Hénon dynamics.
+The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value indicates that this is an inadequate parametric model family.
+Figure 2: Mean $\log\det\Sigma_{\mathrm{model\ error}}$ and 1-standard deviation interval over time for the parametric model family (18) and the Hénon dynamics.
+The fact that $\log\det\Sigma_{\mathrm{model\ error}}$ plateaus higher than its initial value indicates that this is an inadequate parametric model family.
+(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
 Figure 3: Results for the linear dynamics.
-(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
-(b) Inputs as a function of time for the various methods for a specific initialization of $\Sigma$.
-(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
+(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
 Figure 4: Results for the Hénon dynamics.
-(a) Posterior model parameter estimate in proposed algorithm as it evolves over time for a specific initialization of $\Sigma$.
-(b) State of dynamics as a function of time for the various methods for a specific initialization of $\Sigma$.
-(c) Mean model parameter estimate error and 1-standard deviation interval for each method. The randomness is over the initial values of $\Sigma$.
+(a) Model parameter estimate (point value in red and posterior covariance (9) in blue) as it evolves over time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(b) Inputs as a function of time for specific $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
+(c) Mean model parameter estimate error and 1-standard deviation interval. The randomness is over $\theta_{\mathrm{prior}}$, $\Sigma_{\mathrm{prior}}$ and initial $\Sigma$.
 Figure 5: Results for the unicycle dynamics.
-Method/Experiment Linear Dynamics Hénon Dynamics Unicycle Dynamics
-proposed $\mathbf{-77.2}$ $\mathbf{-51.2}$ $\mathbf{-43.5}$
-baseline $-12.1$ $-31.8$ $-7.5$
-TABLE I: $\log\det$ of the $\Sigma_{\mathrm{posterior}}$ at the final timestep for the various dynamics studied.
 VI Conclusion
@@
 It extends to general nonlinear system identification through the Taylor expansion.
-A key module in our algorithm is the online calibration of the noise covariance that greatly impacts performance and allows us to judge the suitability of a parametric model family for the underlying system.
-Our algorithm also has limitations.
+A key module in our algorithm is the online calibration of the noise covariance that also allows us to judge the suitability of a parametric model family for the underlying system.
+Our algorithm however has limitations.
 The use of the trust region, although necessary to ensure the accuracy of the Taylor series, might forbid large changes in the parameters when these are implied by the data.
+The calibration of the covariance is dependent on the current estimate value, which can slow down performance if the current estimate is not good.
 Future work will apply our method to high-dimensional problems and look into theoretical guarantees.
@@
 Toyota Research Institute (TRI) provided funds to assist the authors with their research, but this article solely reflects the opinions and conclusions of its authors and not TRI or any other Toyota entity.
-The NASA University Leadership Initiative (grant ##\##80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity.
+The NASA University Leadership Initiative (grant #\#80NSSC20M0163) provided funds to assist the first author with their research, but this article solely reflects the opinions and conclusions of its authors and not any NASA entity.
 For the first author, this work was also partially funded through the Alexander S.